Titanium is a transition metal used in many alloys because it is extremely strong and lightweight. Titanium tetrachloride (TiCl4) is extracted from titanium oxide (TiO2) using chlorine and coke (carbon). What is the mass of Cl2 gas is needed to react with 1.25 mol of TiO2?



1
Expert's answer
2021-10-10T10:27:32-0400

The reaction is represented by the following equation:

TiO2 + 2Cl2 + C --> TiCl4 + CO2

Therefore,

"m(Cl_2)=1.25mol(TiO_2)\times\frac{2mol(Cl_2)}{1mol(TiO_2)}\times\frac{70.91g(Cl_2)}{1mol(Cl_2)}=177g"


Answer: 177 g